\begin{thebibliography}{998}

%AAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAA
\bibitem{AN1952}Ankeny, N. C. \textit{\color{red}Representations of primes by quadratic forms.} Amer. J. Math. 74 (1952), 913--919. 
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0049931}{MR0049931}.

\bibitem{FG2018}Aletheia-Zomlefer, S. L.; Fukshansky, L.; Garcia, S. R. \textit{\color{red}The Bateman-Horn Conjecture: Heuristics, History, and Applications.} Expo. Math. 38 (2020), no. 4, 430--479. \href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR4177951}{MR4177951}.


%\bibitem{AP1998} Apostol, Tom M. \textit{\color{red}Introduction to analytic number theory.} Undergraduate Texts in Mathematics. Springer-Verlag, New York-Heidelberg, 1976.





%BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBB



%\bibitem{BA1985} Balog, A. \textit{\color{red}On the distribution of $p^{\theta}\\bmod 1$}. Acta. Math. Hungar., 45 (1985), pp. 179--199. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0779531}{MR0779531}.


%\bibitem{BA1983} Balog, A. \textit{\color{red}On the fractional part of $p^{\theta}$}. Arch. Math. (Basel) 40 (1983), no. 5, 434--440. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0707732}{MR0707732}.

%\bibitem{BI1989} Bombieri, E.; Friedlander, J. B.; Iwaniec, H. \textit{\color{red}Primes in arithmetic progressions to large moduli. III.} J. Amer. Math. Soc. 2 (1989), no. 2, 215--224. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0976723}{MR0976723}.


%\bibitem{BP2001} Baker, R. C.; Harman, G.; Pintz, J. \textit{\color{red}The difference between consecutive primes. II.} Proc. London Math. Soc. (3) 83 (2001), no. 3, 532--562. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1851081}{MR1851081}.


\bibitem{BH1962} Bateman, P. T.; Horn, R. A. \textit{\color{red}A heuristic asymptotic formula concerning the distribution of prime numbers.} Math. Comput., 16 (1962), pp. 363--367.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0148632}{MR0148632}.

%\bibitem{BK1985}R.C. Baker, R. C.; Kolesnik, G. \textit{\color{red}On the distribution of $p^{\alpha}$ modulo one.} J. Reine Angew. Math., 356 (1985), pp. 174--193. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0779381}{MR0779381}.



%\bibitem{BO15} Bordelles, Olivier. \textit{\color{red}Some explicit estimates for the M�bius function.} J. Integer Seq. 18 (2015), no. 11, Article 15.11.1, 13 pp.

\bibitem{BZ2007} Baier, S.; Zhao, L. \textit{\color{red}On Primes Represented by Quadratic Polynomials.} CRM Proc. Lecture Notes, 46
American Mathematical Society, Providence, RI, 2008, 159--166.
\href{http://arxiv.org/abs/math/0703284}{Arxiv.org/abs/math/0703284}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2437972}{MR2437972}.

\bibitem{BZ2007B} Baier, S.; Zhao, L. \textit{\color{red}Primes in quadratic progressions on average}. Math. Ann. 338 (2007), no. 4, 963--982. 
\href{http://arxiv.org/abs/math/0605563}{Arxiv.org/abs/math/0605563}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2317757}{MR2317757}.

%\bibitem{BS09} Banks, William D.; Shparlinski, Igor E. \textit{\color{red}Prime numbers with Beatty sequences}. Colloq. Math.  115  (2009),  no. 2, %147-157. 
%\bibitem{BR95} Baker, R. C.; Harman, G.; Rivat, J. \textit{\color{red}Primes of the form $n^c$.} J. Number Theory  50  (1995),  no. 2, 261-277. 
%\bibitem{CC00} Chris K. Caldwell. \textit{\color{red}An Amazing Prime Heuristic, Preprint, 2000.

%CCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCC

%\bibitem{CA2004}Cojocaru, Alina Carmen. \textit{\color{red}Questions about the reductions modulo primes of an elliptic curve}. Number theory, 61--79, CRM Proc. Lecture Notes, 36, Amer. Math. Soc., Providence, RI, 2004.

\bibitem{CJ2011} Cilleruelo, J. \textit{\color{red} The least common multiple of a quadratic sequence.} Compos. Math. 147 (2011), no. 4, 1129--1150.
%\href{http://arxiv.org/abs/math/0703284}{Arxiv.org/abs/math/0703284}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2822864}{MR2822864}.


\bibitem{CM1993} Coleman, M. D. \textit{\color{red}The Rosser-Iwaniec sieve in number fields, with an application.} Acta Arith. 65 (1993), no. 1, 53--83. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1239243}{MR1239243}.

%DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

\bibitem{DI1982} Deshouillers, J. M.; Iwaniec, H. \textit{\color{red}On the greatest prime factor of $n^2 + 1$.} Ann. Inst. Fourier (Grenoble) 32 (1982), no. 4, 1--11.
%\href{http://arxiv.org/abs/math/0703284}{Arxiv.org/abs/math/0703284}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0694125}{MR0694125}. 

\bibitem{DH2000} Davenport, H. \textit{\color{red}Multiplicative number theory.} Second Edition. Graduate Texts in Mathematics. Springer-Verlag, Berlin, 1980. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=1790423}{MR1790423}.
	
\bibitem{DLMF} NIST \textit{\color{red}Digital Library of Mathematical Functions}. F. W. J. Olver, ..., and M. A. McClain, eds.; 2019. \href{http://dlmf.nist.gov/}{dlmf.nist.gov}.




%EEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE

\bibitem{EL1985}  Ellison, William; Ellison, Fern. \textit{\color{red}Prime numbers}. A Wiley-Interscience Publication. John Wiley and Sons, Inc., New York; Hermann, Paris, 1985. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0814687}{MR0814687}.


\bibitem{EL1900} Euler, Leonhard.  \textit{\color{red}De Numeris Primis Valde Magnis}, Novi Commentarii academiae scientiarum Petropolitanae 9, 1764, pp. 99--153. \href{http://eulerarchive.maa.org/docs/originals/E283.pdf}{eulerarchive.maa.org/docs/originals/E283}.

\bibitem{EH2008}Everest, G.; Harman, G. \textit{\color{red}On primitive divisors of $n^2+b.$} 
London Math. Soc. Lecture Note Ser., 352
Cambridge University Press, Cambridge, 2008, 142--154.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2428520}{MR2428520}.

\bibitem{EP1952}Erdos, P. \textit{\color{red}On the greatest prime factor of  $\prod_{1\leq k\leq x} f(k)$.} J. Lond. Math. Soc. 27, 379--384 (1952).
\href{https://zbmath.org/0046.04102}{zbMath0046.04102}.




%\bibitem{EP1960} Erdos, P.  \textit{\color{red}On a problem of G. Golomb}. J. Austral. Math. Soc. 2 1961/1962 1-8. 

%FFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFF

\bibitem{FI2022} Friedlander, J.; Iwaniec, H. \textit{\color{red}Coordinate distribution of Gaussian primes}. J. Eur. Math. Soc. 24 (2022), 737--772. \href{http://arxiv.org/abs/1811.05507}{Arxiv.org/abs/1811.05507}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR4397031}{MR4397031}.


\bibitem{FI2010} Friedlander, J.; Iwaniec, H. \textit{\color{red}Opera de cribro}. American Mathematical Society Colloquium Publications, 57. AMS, Providence, RI, 2010.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2647984}{MR2647984}.

\bibitem{FM1988} Filaseta, M. \textit{\color{red}Prime values of irreducible polynomials.} Acta Arith. 50 (1988), no. 2, 133--145. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0945264}{MR0945264}.

%GGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGG

\bibitem{GL2010} Goldoni, L. \textit{\color{red}Prime Numbers And Polynomials}, PhD Thesis, Universita` Degli Studi Di Trento, 2010.

%\bibitem{GS1986} Gritsenko, S. A. \textit{\color{red}On a problem of I. M. Vinogradov.} Mat. Zametki, 39 (1986), pp. 625--650. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0850799}{MR0850799}.


\bibitem{GM2000} Granville, A.; Mollin, R. A. \textit{\color{red}Rabinowitsch revisited}. Acta Arith. 96 (2000), no. 2, 139--153. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1814449}{MR1814449}.


	
%	\bibitem{BR19} Roger Baker. \textit{\color{red}A theorem of Bombieri-Vinogradov type with few exceptional moduli,} arXiv:1905.12488.
	
	
%	\bibitem{BT19}Oleksiy Klurman, Alexander P. Mangerel, Joni Teravainen. \textit{\color{red}Multiplicative functions in short arithmetic progressions,} arXiv:1909.12280.

%HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH							

\bibitem{HC1967} Hooley, C. \textit{\color{red} On the greatest prime factor of a quadratic polynomial.} Acta Math., 117:281-299, 1967.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=0204383}{MR0204383}.

\bibitem{HG2008}Harman, G. \textit{\color{red}On values of $ n^2+1$  free of large prime factors.} Arch. Math. (Basel) 90 (2008), no. 3, 239--245.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2391358}{MR2391358}.

\bibitem{HG2024}Harman, G. \textit{\color{red}On Two problems on the greatest prime factor of $n^2 + 1$.} Acta Arithmetica Vol. 213 (2024), 273--287.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2391358}{MR2391358}.
							
\bibitem{HG2007} Harman, G. \textit{\color{red}Prime-detecting sieves}. London Mathematical Society Monographs Series, 33. Princeton University Press, Princeton, NJ, 2007.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2331072}{MR2331072}.

\bibitem{HL2001} Harman, G.;  Lewis, P. \textit{\color{red}Gaussian primes in narrow sectors.} Mathematika 48 (2001), 119--135.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1996365}{MR1996365}.

%\bibitem{HG1991} Harman, G. \textit{\color{red}Fractional and integral parts of $p^{\lambda}$.} Acta Arith. 58 (1991), 141--152. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1121075}{MR1121075}.

%\bibitem{HG1983} Harman, G. \textit{\color{red}On the distribution of $\{\sqrt{p}\}$ modulo one}. Mathematika  30  (1983),  no. 1, 104--116. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0720954}{MR0720954}.

%\bibitem{HR1995} Harman, Glyn; Rivat, Joel. \textit{\color{red}Primes of the form $[p^c ]$ and related questions.} Glasgow Math. J.  37  (1995),  no. 2, %131-141.

%\bibitem{HG2016} Harman, Glyn. \textit{\color{red}Primes in Beatty sequences in short intervals}. Mathematika  62  (2016),  no. 2, 572-586. 

\bibitem{HL1923} Hardy, G. H. Littlewood J.E. \textit{\color{red}Some problems of Partitio numerorum III: On the expression of a number as a sum of primes}. Acta Math. 44 (1923), No. 1, 1--70.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1555183}{MR1555183}.

%\bibitem{HR59} Hardy, G. H. Ramanujan: \textit{\color{red}Twelve lectures on subjects suggested by his life and work.}  Chelsea Publishing Company, New York 1959.



%\bibitem{HG2016}Hurst, Greg. \textit{\color{red}Computations of the Mertens function and improved bounds on the Mertens conjecture}. Math. Comp. 87 (2018), no. 310, 1013-1028 or arxiv.1610.08551.



%\bibitem{HW08} Hardy, G. H.; Wright, E. M. An introduction to the theory of numbers. Sixth edition. Revised by D. R. Heath-Brown and J. H. Silverman. %With a foreword by Andrew Wiles. Oxford University Press, Oxford, 2008. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2445243}{MR2445243}.
%IIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIII

\bibitem{IK2004} Iwaniec, H.; Kowalski, E. \textit{\color{red}Analytic number theory.} Amer. Math. Soc. Colloquium Publications, 53. American Mathematical Society, Providence, RI, 2004.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=2061214}{MR2061214}.	

\bibitem{IH1978} Iwaniec, H. \textit{\color{red}Almost-primes represented by quadratic polynomials.} Invent. Math. 47 (1978), no. 2, 171--188. 
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0342464}{MR0342464}.


%JJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJJ

%\bibitem{JW2003} Jacobson, M. J.; Williams, Hugh C. \textit{\color{red}New quadratic polynomials with high densities of prime values}. Math. Comp. 72 (2003), no. 241, 499--519.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1933834}{MR1933834}.

%KKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKK															
%\bibitem{KA99} Kumchev, A. \textit{\color{red}On the distribution of prime numbers of the form $[n^c ]$} . Glasg. Math. J.  41  (1999),  no. 1, 85-%102.

%\bibitem{KE05} Emmanuel Kowalski. \textit{\color{red}Analytic problems for elliptic curves}. arXiv:math/0510197.
																
%\bibitem{KJ92} Keiper, J. B. \textit{\color{red}Power series expansions of Riemann's $\xi$ function. Math. Comp. 58 (1992), no. 198, 765-773.

%\bibitem{KL12} Kunik, Matthias; Lucht, Lutz G. \textit{\color{red}Power series with the von Mangoldt function. Funct. Approx. Comment. Math. 47 %(2012),  part 1, 15-33. 

%\bibitem{KR1979}Kaufman, R. M. \textit{\color{red}The distribution of $\{\sqrt{p}\}$.} Math. Notes 26 (1979), no. 3-4, 739--743 (1980). \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0552712}{MR0552712}.



%\bibitem{KR09} Jacob Korevaar. \textit{\color{red}Prime pairs and the zeta function, Journal of Approximation Theory 158 (2009) 69-96.

%\bibitem{KT06} Kotnik, Tadej; te Riele, Herman. \textit{\color{red}The Mertens conjecture revisited}. Algorithmic number theory, 156-167, Lecture %Notes in Comput. Sci., 4076, Springer, Berlin, 2006.

%LLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLL

\bibitem{LN1997} Lidl, R.; Niederreiter, H. \textit{\color{red}Finite fields.} Second edition. Encyclopedia of Mathematics and its Applications, 20. Cambridge University Press, Cambridge, 1997. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1429394}{MR1429394}.

%\bibitem{LL95} Lucht, Lutz. \textit{\color{red}Ramanujan expansions revisited. Arch. Math. (Basel) 64 (1995), no. 2, 121-128.

%\bibitem{LL10} Lucht, Lutz G. \textit{\color{red}A survey of Ramanujan expansions. Int. J. Number Theory 6 (2010), no. 8, 1785-1799. 

\bibitem{LR2012} Lemke Oliver, R. J. \textit{\color{red}Almost-primes represented by quadratic polynomials.} Acta Arith. 151 (2012), no. 3, 241--261.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2860953}{MR2860953}.

%MMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMM

\bibitem{MA2009} Matomaki, Kaisa. \textit{\color{red}A note on primes of the form $p=aq^2+1$}. Acta Arith. 137 (2009), no. 2, 133--137. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2491532}{MR2491532}.

\bibitem{MJ2023} Merikoski, J. \textit{\color{red} On the largest prime factor of  $n^2 + 1$}. J. Eur. Math. Soc. (JEMS) 25 (2023), no. 4, 1253-1284.
\href{http://arxiv.org/abs/1908.08816}{arxiv.org/abs/1908.08816}, 
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=4577964}{MR4577964}.


\bibitem{MK1986} McCurley, K. S. \textit{\color{red}The smallest prime value of $x^n +a$} . Canad. J. Math.  38  (1986),  no. 4, 925--936.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0840616}{MR0840616}.

%\bibitem{MM13} Mariusz Mirek. \textit{\color{red}Roth Theorem in the Piatetski-Shapiro primes}, arXiv:1305.0043. 

%\bibitem{MR19} James Maynard, Zeev Rudnick. \textit{\color{red}A lower bound on the LCM of polynomial sequences}, arXiv:1910.13218.

%\bibitem{MK1984} McCurley, Kevin S. \textit{\color{red}Prime values of polynomials and irreducibility testing}. Bull. Amer. Math. Soc. (N.S.) 11 (1984), no. 1, 155-158. 

%\bibitem{MS07} Maier, H.; Stewart, C. L. \textit{\color{red}On intervals with few prime numbers. J. Reine Angew. Math. 608 (2007), 183-199.

\bibitem{MV2007} Montgomery, H. L., Vaughan, R. C. \textit{\color{red}Multiplicative number theory. I}. Classical theory. Cambridge
University Press, Cambridge, 2007. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=2378655}{MR2378655}.

%NNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNN

\bibitem{NT1921}Nagel, T. \textit{\color{red} 
	Generalisation d'un theoreme de Tchebycheff.} Journ. de Math. (8) 4, 343--356 (1921).
\href{https://zbmath.org/48.1173.01}{zbMath48.1173.01}.

\bibitem{NW2000} Narkiewicz, W. \textit{\color{red}The development of prime number theory. From Euclid to Hardy and Littlewood}. Springer Monographs in Mathematics. Springer-Verlag, Berlin, 2000. 
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1756780}{MR1756780}.

%PPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP

%\bibitem{PF2010} Pillichshammer, F. \textit{\color{red} Euler's constant and averages of fractional parts}. Amer. Math. Monthly 117 (2010), no. 1, 78--83. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2541431}{MR2541431??}.

\bibitem{PJ2009} Pintz, Janos. \textit{\color{red}Landau's problems on primes}. J. Theory. Nombres Bordeaux 21 (2009), no. 2, 357--404.
%\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2541431}{MR2541431}.


%\bibitem{PK1957} Prachar, Karl. \textit{\color{red} Handbuch der Lehre von der Verteilung der Primazahlen [1909]} Berlin, Springer, 1957. 

%\bibitem{RM13} Ram Murty, M. \textit{\color{red}Ramanujan series for arithmetical functions. Hardy-Ramanujan J.  36  (2013), 21-33.

%RRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRR

%\bibitem{RD1996} Redmond, Don.  \textit{\color{red}Number theory. An introduction}. Monographs and Textbooks in Pure and Applied Mathematics, 201. Marcel Dekker, Inc., New York, 1996. %\href{http://arxiv.org/abs/1807.08899}{Arxiv.org/abs/1807.08899}. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3434886}{MR3434886???}.

\bibitem{RI2015} Rivin, I. \textit{\color{red}Some experiments on Bateman-Horn.}
\href{http://arxiv.org/abs/1508.07821}{Arxiv.org/abs/1508.07821}.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3434886}{MR3434886???}.

\bibitem{RP1996} Ribenboim, P. \textit{\color{red}The new book of prime number records}, Berlin, New York: Springer-Verlag, 1996.
%\href{http://arxiv.org/abs/1508.07821}{Arxiv.org/abs/1508.07821}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1377060}{MR1377060}.

%\bibitem{RS01} Rivat, Joel; Sargos, Patrick. \textit{\color{red}Nombres premiers de la forme $[n^ c]$.} Canad. J. Math.  53  (2001),  no. 2, 414-433.

%SSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSS
																						
%\bibitem{SJ08} Jorn Steuding, Marc Technau. \textit{\color{red}The Least Prime Number in a Beatty Sequence,} arXiv:1512.08382.

%\bibitem{SD1964} Shanks, D. \textit{\color{red}An analytic criterion for the existence of infinitely many primes of the form  $(n^2+1)/2$.} Illinois J. Math. 8 (1964), 377--379. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0166145}{MR0166145}.

%\bibitem{SD1960} Shanks, D. \textit{\color{red}On numbers of the form $n^4 +1$ }. Math. Comput.  15,  1961, 186--189.
%\href{http://arxiv.org/abs/1508.07821}{Arxiv.org/abs/1508.07821}. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0120184}{MR0120184}.

\bibitem{SD1959} Shanks, D. \textit{\color{red}A sieve method for factoring numbers of the form $n^2 +1$ }. Math. Tables Aids Comput.  13,  1959, 78--86.
%\href{http://arxiv.org/abs/1508.07821}{Arxiv.org/abs/1508.07821}.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0105784}{MR0105784}.



%\bibitem{SP80} Spilker, Jurgen. \textit{\color{red}Ramanujan expansions of bounded arithmetic functions. Arch. Math. (Basel) 35 (1980), no. 5, 451-%453.
																							
																							
%\bibitem{SS94} Schwarz, Wolfgang; Spilker, Jurgen. \textit{\color{red}Arithmetical functions. An introduction to elementary and analytic properties of %arithmetic functionss. London Mathematical Society Lecture Note Series, 184. Cambridge University Press, Cambridge, 1994.

%TTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTT	
%\bibitem{TG15} Tenenbaum, Gerald. \textit{\color{red}Introduction to analytic and probabilistic number theory. Translated from the Third French %edition. American Mathematical Society, Rhode Island, 2015.

%\bibitem{TA2000} Toth, Arpad. \textit{\color{red}Roots of quadratic congruences}. Internat. Math. Res. Notices  2000,  no. 14, 719--739. 
%\href{http://arxiv.org/abs/1508.07821}{Arxiv.org/abs/1508.07821}. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1776618}{MR1776618}.
																							
%VVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVV
%\bibitem{VR97} Vaughan, R. C. \textit{\color{red}The Hardy-Littlewood method. Second edition. Cambridge Tracts in Mathematics, 125. Math. Soc. (N.S.) %44 (2007), no. 1, 1-18. Cambridge University Press, Cambridge, 1997. 

%\bibitem{VR73} Vaughan, R. C. \textit{\color{red}Some applications of Montgomery's sieve}. J. Number Theory 5 (1973), 64-79.

%\bibitem{VI1976}Vinogradov, I. M. \textit{\color{red}Special Variants of the Method of Trigonometric Sums.} Izda. Nauka, Moskow 1976. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0469878}{MR0469878}.


%WWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWW																							
%\bibitem{WN66} Wiener, Norbert. \textit{\color{red}Generalized harmonic analysis and Tauberian theorems. M.I.T. Press, Cambridge, Mass., London 1966.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage[top=3.5cm, bottom=3.5cm, left=3cm, right=3cm]{geometry}
\usepackage[latin1]{inputenc}
\usepackage{csquotes}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphics}
\usepackage{float}
\usepackage[hang,flushmargin]{footmisc} 

\usepackage{amsmath,amsthm,amssymb,graphicx, multicol, array}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{xcolor}

\usepackage[pagebackref,bookmarks,colorlinks,breaklinks]{hyperref}
\hypersetup{linkcolor=blue,citecolor=red,filecolor=blue,urlcolor=blue} 

\usepackage{tabto}

\numberwithin{equation}{section}

\usepackage{tikz}
\usepackage{pgfplots}

\usetikzlibrary{matrix}

\usepackage{mathrsfs}

\DeclareMathOperator{\lcm}{lcm}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]

\newtheorem{conj}{Conjecture}[section]

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\Q}{\mathbb{Q}}

\newcommand{\C}{\mathbb{C}}
\newcommand{\F}{\mathbb{F}}

\usepackage{fancyhdr}
\title{Prime Values of the Euler Polynomial}
\date{}
\author{N. A. Carella}

\begin{document}

\thispagestyle{empty}
\date{}
\maketitle

\begin{abstract}
This note proposes an effective lower bound for the number of  primes in the quadratic progression $p=n^2+1 \leq x$ as $x \to \infty$.\let\thefootnote\relax\footnote{ \today \date{} \\
	\textit{AMS MSC}: Primary 11N322, Secondary 11N05. \\
	\textit{Keywords}: Prime number; Polynomial prime value.}
\end{abstract}


\section{Introduction} \label{S9090}\hypertarget{S9090}
As early as 1760, Euler was developing the theory of prime values of polynomials. In fact, Euler computed a very large table of the primes $p=n^2+1$, see {\color{red}\cite[p.\ 123]{EL1900}}. Likely, the prime values of polynomials was studied by other researchers before Euler. Later, circa 1910, Landau posed an updated question of the same problem about the primes values of the polynomial $n^2+1$. A fully developed conjecture, based on circle methods, was demonstrated about two decades later.

\begin{conj} \label{conj9090.112}\hypertarget{conj9090.112}{\normalfont ({\color{red}\cite[Conjecture E, p.\,46]{HL1923}})} Let $x \gg 1$ be a large number, let $\Lambda(n)$ be the vonMangoldt function, and let $\chi(n)=(n\mid p)$ be the quadratic symbol modulo $p$. Then
\begin{equation} \label{eq9090.100}
\sum_{n\leq x} \Lambda \left( n^2+1 \right )
=a_2x+O\left (\frac{ x}{\log x }\right ),
\end{equation}
where the density constant
\begin{equation} \label{eq9090.120}
a_2=\prod_{p \geq 3}\left( 1-\frac{\chi(-1)}{p-1} \right )=1.37281346 \ldots .
\end{equation}
\end{conj}

The claim in Conjecture E is specific to the polynomial $f(t)=t^2+1$ and the next conjecture F in \cite{HL1923} covers the general admissible quadratic polynomial $f(t)=at^2+bt+c\in \Z[t]$. An \textit{admissible} polynomial $f(t)\in \Z[t]$ has a relatively prime image $\gcd(f(\Z))=1$, see {\color{red}\cite[p.\;385]{FI2010}}, \cite{FM1988} and similar sources for detailed explanations. A survey of the subsequent developments appears in {\color{red}\cite[p.\ 342]{NW2000}}, {\color{red}\cite[Section 19]{PJ2009}}, \cite{BZ2007}, and similar references. Other related problems are discussed in \cite{BH1962}, \cite{BZ2007}, \cite{GL2010}, \cite{FI2010}, {\color{red}\cite[p. 343]{NW2000}}, {\color{red}\cite[p. 405]{RP1996}}, \cite{PJ2009}. \\

There are many approaches and techniques toward the proof of \hyperlink{conj9090.112}{Conjecture} \ref{conj9090.112} on the prime values of the admissible irreducible polynomial $n^2+1$ and other admissible and irreducible polynomials. Two of the leading techniques are essentially approximations to the actual problem: 
\begin{itemize}
\item Determining the \textit{largest prime factor }of the product $\prod_{n\leq x}(n^2+1)$. \\
\item Determining the \textit{smallest coefficient} $m$ in the representation $p=n^2+m^2$. \\
\end{itemize}

There is a large literature dedicated to the determination of the largest prime factors of admissible irreducible polynomials. The analysis of these results are based on the Chebyshev-Hooley method, which is a difficult analytical technique, see \cite{HC1967}, \cite{DI1982}, \cite{MJ2023}, et alii.  In addition, the properties of the primitive divisors of pure polynomials such as  $f(t)=t^2+c$ and some related topics are studied in \cite{EH2008}, \cite{HG2008}, \cite{HG2024}, et alii. An abridged table of the known results focused on the polynomial $f(t)=t^2+1$ are listed in \autoref{table9090-A}. Here, the symbol $P(n)\geq1$ denotes the largest prime factor of an integer $n\geq1$.

\vskip .1 in 
\begin{table}[H]
	\setlength{\tabcolsep}{0.5cm}
	\renewcommand{\arraystretch}{1.750092}
	\setlength{\arrayrulewidth}{0.80pt}
	\centering
	\begin{tabular}{l|l}
		
		Estimate of the largest prime divisor	& Author and reference \\
		\hline
		$P(n^2+1)/n\to \infty $ as $ n\to\infty$	&  Chebyshev, see \cite{HC1967}.\\
		\hline
		$P(n^2+1)\geq n(\log n)^{\varepsilon}$ as $ n\to\infty$	& Nagell, see \cite{NT1921}. \\
		\hline
		$P(n^2+1)\geq n(\log n)^{c\log\log\log n}$ as $ n\to\infty$		&  Erdos, see \cite{EP1952}.\\
		\hline
		$P(n^2+1)\geq n^{11/10}$ as $ n\to\infty$		&  Hooley, see \cite{HC1967}.\\
		\hline
		$P(n^2+1)\geq n^{1.202}$ as $ n\to\infty$		&  Iwaniec, see \cite{DI1982}.\\
		\hline
		$P(n^2+1)\geq n^{1.279}$ as $ n\to\infty$		&  Merikoski, see \cite{MJ2023}.\\
	\end{tabular}  
	\vskip .1 in		
	\caption{Estimated magnitude of prime divisor of $n^2+1$}
	\label{table9090-A}
\end{table}

Likewise, there is a large literature dedicated to the determination of smallest coefficient in the representation $p=n^2+m^2$. In terms of complex gaussian numbers $\pi=n+im\in\Z[i]$, the rational integer $n^2+m^2=\pi \overline {\pi}$ is a prime number infinitely often if and only if the complex line $\Im  m \;\pi=1\subset \C$ contains infinitely many gaussian primes $\pi=n+i$. Moreover, conditional on the generalized RH, there are a few proofs for the existence of infinitely many gaussian primes $\pi=a+ib\in \Z[i]$ on the complex strip $0\leq \Im  m \;\pi\ll \log p\subset \C$, confer {\color{red}\cite[Theorem III]{AN1952}}, et alii.  Unconditionally, there are various results regarding the existence of infinitely many gaussian primes $\pi=a+ib\in \Z[i]$ on the complex strip $0\leq \Im  m \;\pi\ll p^{\alpha}\subset \C$, where $\alpha<1/2$. More generally, the proof for quadratic forms $p=Q((m,n)=am^2+bmn+cm^2$, with $m\ll p^{\alpha}$, and a survey of previous results appears in \cite{CM1993}. These results are equivalent to the results for complex sectors $\alpha\leq \arg \pi \leq \beta\subset \C$. Some topics in the advanced algebraic and analytic theory of the gaussian number field $\Q(i)$ and primes as sums of two squares, such as the primes $p=n^2+1$, are treated in many papers, see \cite{FI2022}, {\color{red}\cite[Chapter 12]{HG2007}} et alii. An abridged table of the progress is listed on \autoref{table9090-B}.  
\vskip .1 in 
\begin{table}[H]
	\setlength{\tabcolsep}{0.5cm}
	\renewcommand{\arraystretch}{1.750092}
	\setlength{\arrayrulewidth}{0.80pt}
	\centering
	\begin{tabular}{l|l|l}
		Smallest coefficient $m$	& Criterion&Authors and references \\
		\hline
		$m\ll\log  p$	&Conditional  &Ankeny 1952, \cite{AN1952}\\		
		\hline
		$m\ll p^{1/4}$	& Unconditional& Simple exercise.\\
		\hline
		$m\ll p^{0.1631}$	&Unconditional  &Coleman 1993, \cite{CM1993}\\
		
		\hline
		$m\ll p^{0.119}$	&Unconditional  &Harman-Lewis 2001, \cite{HL2001}\\
	\end{tabular}
	\vskip .1 in		
	\caption{The prime $p=n^2+m^2$ and the smallest $m$}
	\label{table9090-B}
\end{table}

Some partial results are proved in \cite{GM2000}, \cite{MA2009}, \cite{BZ2007}, \cite{DI1982}, \cite{IH1978}, \cite{LR2012}, et alii. The smallest expected prime value of certain admissible polynomials $x^d+a\in\Z[x]$ is studied in \cite{MK1986}. Assuming the Elliott-Halberstam conjecture, there is a discussion in {\color{red}\cite[p.\ 5]{BZ2007}} concerning the existence of infinitely many primes of the form $p=an^2+1$, with $a =O(p^{\varepsilon})$, and $\varepsilon>0$. Extensive discussion on this later conjecture and the level of distribution of the moduli are given in {\color{red}\cite[p.\ 406]{FI2010}}. The results for the associated least common multiple problem $\log \lcm [f(1)f(2)\cdots f(n)]$ appears in \cite{CJ2011}, and the recent literature. This note proposes the following partial result.

\begin{thm} \label{thm9090.119}\hypertarget{thm9090.119} Let $x \geq 1$ be a large number. Then, 
\begin{equation} \label{eq909.19}
 \sum_{\sqrt{x}< n\leq \sqrt{2x}} \Lambda(n^2+1)\gg \sqrt{x}\left( 1+O\left(e^{-c\sqrt{\log x}} \right) \right).
\end{equation}
\end{thm}

This result seems to resolve the asymptotic part of the polynomial prime values problem for the admissible polynomial $f(x)=x^2+1$. The proof, based on standard elementary methods in number theory, appears in \hyperlink{S2099}{Section} \ref{S2099}. The supporting materials are developed in \hyperlink{S2070}{Section} \ref{S2070}, \hyperlink{S2099MT}{Section} \ref{S2099MT} and \hyperlink{S2099ET}{Section} \ref{S2099ET}. \\

Some heuristics, and discussions on the difficulty and complexity of estimating or computing the constant appear in \cite{BH1962}, {\color{red}\cite[Section 3.3]{FG2018}}, \cite{RI2015}, and other papers on the Bateman-Horn Conjecture.

\section{Linear to Quadratic Identity} \label{S2070}\hypertarget{S2070}
Among the representations of the characteristic function of square integers $n \in [1,x]$ there is the double finite sum
\begin{equation}\label{eq2070.130d}
\mathcal{I}_2(n)=\frac{1}{N}\sum_{m \leq \sqrt{x},}\sum_{0\leq u<N} e^{i2\pi \left(m^2-n \right)u/N } =
\begin{cases}
1&\text{  if }  n=m^2,\\
0&\text{  if }  n\ne m^2,\\
\end{cases} 
\end{equation}
where $N>1$ is a parameter and $x>1$ is a real number. The weighted characteristic function of prime number, better known as the vonMangoldt function, is defined by

\begin{equation}\label{eq2070.130f}
\Lambda(n)=
	\begin{cases}
		\log p &\text{  if }  n=p^v,\\
		0&\text{  if }  n\ne p^v,\\
	\end{cases} 
\end{equation}
where $n=p^v$ is a prime power.\\

Merging these two characteristic functions produces the
\textit{quadratic to linear identity}.

\begin{lem}\label{lem2070.130}\hypertarget{lem2070.130} If $n\geq$ is an integer, then
\begin{equation}\label{eq2070.130i}
\mathcal{I}_2(n)\Lambda(n+1) =
\begin{cases}
\Lambda(m^2+1)&\text{  if }  n=m^2 \text{ is a square,}\\
0&\text{  if }  n\ne m^2 \text{ is not a square.}\\
\end{cases} 
\end{equation}
\end{lem}  

The application of the characteristic function \eqref{eq2070.130d} presented here uses the choice of parameter $N\approx x$ or $N\sim x$ and $N$ prime. For a fixed $n\in[1,x]$, this ensures that the congruence equation 
\begin{equation} \label{eq2070.100b}
	z^2-n\equiv 0\bmod N ,
\end{equation}
where $z\leq \sqrt{x}$, and the integer equation 
\begin{equation} \label{eq2070.100c}
	z^2-n= 0
\end{equation} 
have exactly the same integer solution as specified in \eqref{eq2070.130d}. The matched  parameters ensures that the double exponential sum \eqref{eq2070.130d} works as a characteristic function of square integers $n\leq x$ as specified. Mismatched parameters produce \textit{quasi characteristic functions}. For example,  
\begin{equation}\label{eq2070.130g}
\frac{1}{N}\sum_{m \leq x^{3/4},}\sum_{0\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }
\end{equation}
is not a characteristic function of the squared integers for large $N\approx x$.

\section{Fibers and Multiplicities} \label{S2095SS} \hypertarget{S2099SS}
The cardinalities and multiplicities of the fibers occurring in the estimate of the error term are computed in this section. The notation $[x]=x-{x}$ denotes the integer value  of the quantities $x>0$.
\begin{lem}  \label{lem2099SS.300S}\hypertarget{lem2099SS.300S} Let $N$ be a prime and define the maps
	\begin{equation}\label{eq2099SS.300-m}
		\alpha(m,n)\equiv (m^{2}-n)\bmod N\quad \text{ and } \quad 
		\beta(u,v)\equiv uv\bmod N.
	\end{equation}	
	If $\Lambda(n^2+1)=0$ for all $n\in (N,2N)$, then the fibers $\alpha^{-1}(r)$ and $\beta^{-1}(r)$ of an element $0\ne r\in \F_N$
	have the cardinalities 
	\begin{equation}\label{eq2099SS.300-f}
		\#	\alpha^{-1}(r)\leq [\sqrt{2N}]-[\sqrt{N}]-1\quad \text{ and }\quad \#\beta^{-1}(r)=	[\sqrt{N}]
	\end{equation}
	respectively.
\end{lem}
\begin{proof}[\textbf{Proof}] Given a fixed $m\in \mathscr{M}=\{\sqrt{N}<m\leq \sqrt{2N}\}$, the map 
	\begin{equation}\label{eq9900Q.300-m1}						  		
		\alpha: \mathscr{M}\times[1,N-1]\longrightarrow\F_N\quad  \text{ defined by }\quad  \alpha(m,n)\equiv (m^{2}-n)\bmod N,
	\end{equation}
	where $n\ne m^2$,	is one-to-one but not surjective since the value 
	\begin{equation}\label{eq9900Q.300-m2}						  		
		\alpha(m,n)= (m^{2}-n)\equiv 0 \bmod N
	\end{equation} is missing. This follows from the fact that the map $n\longrightarrow m^2-n \bmod N$ is a permutation the nonezero elements of the finite field $\F_N$,
	see {\color{red}\cite[Chapter 7]{LN1997}} for more details on permutation functions of finite fields. Thus, as $(m,n)\in \mathscr{M}\times[1,N-1]$ varies, a value $0\ne r=\alpha(m,n)\in\F_N$ is repeated at most $[\sqrt{2N}]-[\sqrt{N}]-1$ times since the hypothesis implies that $m^2-n\ne0$ for all $(m,n)$. This verifies that the cardinality of the fiber is
	\begin{eqnarray}\label{eq9900Q.300-f1}
		\#	\alpha^{-1}(r)&=&	\#\{(m,n):r\equiv (m^{2}-n)\bmod N:1\leq m\leq \sqrt{N} \text{ and } n<N/2\}\nonumber\\[.3cm]
		&=& [\sqrt{2N}]-[\sqrt{N}]-1.
	\end{eqnarray}		
	Similarly, given a fixed $u\in [1,\sqrt{N}]$, the map 
	\begin{equation}\label{eq9900Q.300-m3}
		\beta:[1,\sqrt{N}]\times [1,N-1]\longrightarrow\F_N\quad  \text{ defined by }\quad  \beta(u,v)\equiv uv\bmod N,
	\end{equation}
	is one-to-one. Here the map $v\longrightarrow uv \bmod N$ permutes the nonzero elements of the finite field $\F_N$. Thus, as $(u,v)\in [1,\sqrt{N}]\times [1,N-1]$ varies, each value $0\ne r=\beta(u,v)\in\F_N$ is repeated exactly $[\sqrt{N}]$ times. This verifies that the cardinality of the fiber is
	\begin{eqnarray}\label{eq9900Q.300-f2}
		\#	\beta^{-1}(r)&=&	\#\{(u,v):r\equiv uv\bmod N:1\leq u\leq \sqrt{x} \text{ and }1\leq v< N\}\nonumber\\[.3cm]
		&=&[\sqrt{x}].
	\end{eqnarray}
	
	Now each value $r=\alpha(m,n)\ne0$ (of multiplicity at most $[\sqrt{2N}]-[\sqrt{N}]-1$ in $	\alpha^{-1}(r)$), is matched to $r=\alpha(m,n)=\beta(u,v)$ for some $(u,v)$, (of multiplicity exactly $[\sqrt{N}]$ in $	\beta^{-1}(r)$). Comparing \eqref{eq9900Q.300-f1} and \eqref{eq9900Q.300-f2} proves that
	\begin{equation}\label{eq9900Q.300-m4}
		\# \alpha^{-1}(r)=[\sqrt{2N}]-[\sqrt{N}]-1<\# \beta^{-1}(r)=[\sqrt{N}].
	\end{equation}
\end{proof}

\section{Estimate for the Main Term} \label{S2099MT}\hypertarget{S2099MT}
In this analysis the relationship $N\sim x$ between the real number $x$ and the prime $N$ is explicated in \hyperlink{S2070}{Section} \ref{S2070}. 
\begin{lem} \label{lem2099MT.100A}\hypertarget{lem2099MT.100A} If $x \geq 1$ is a large number and $N\sim x$ is a large prime, then
\begin{equation} \label{eq2099.100d}
\frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{x}< m\leq \sqrt{2x}}1\gg \sqrt{x}\left( 1+O\left(e^{-c\sqrt{\log x}} \right) \right),\nonumber
\end{equation}
where $c>0$ is a constant.	
\end{lem}

\begin{proof}[\textbf{Proof}] 
Set the parameter $N\sim x$ as a prime greater than $x$. By hypothesis the subsum
\begin{equation} \label{eq2099.100f}
\sum_{\sqrt{x}< m\leq \sqrt{2x}} \Lambda(m^2+1)=\sum_{x<m^2=n\leq 2x} \Lambda(n+1)=0
\end{equation} 
does not contribute to the main term since there are no primes of the form $p=n^2+1\in (x,2x]$. Thus, the main term can be rewritten and estimated as 
\begin{eqnarray} \label{eq2099.100h}
M(x)&=&\frac{1}{N}\sum_{x< n\leq 2x} \Lambda(n+1)\sum_{\sqrt{x}< m\leq \sqrt{2x}}1\\	[.3cm]
	&=&\frac{(\sqrt{2}-1)\sqrt{x}+O(1)}{N}\sum_{x< n\leq 2x}  \Lambda(n+1)  \nonumber\\[.3cm]
	&=&\frac{(\sqrt{2}-1)\sqrt{x}+O(1)}{N}\left( x+O\left(xe^{-c\sqrt{\log x}} \right) \right) \nonumber\\[.3cm]
		&\gg&\sqrt{x}\left( 1+O\left(e^{-c\sqrt{\log x}} \right) \right) \nonumber,
\end{eqnarray}
where $c>0$ is a constant. The third line in \eqref{eq2099.100h} follows from the prime number theorem, see {\color{red}\cite[27.12.E6]{DLMF}}, \cite{EL1985}, \cite{IK2004}, {\color{red}\cite[Theorem 6.9]{MV2007}}. Lastly, the last line follows from $N\sim x$.
\end{proof}

\section{Estimate for the Error Term} \label{S2099ET} \hypertarget{S2099ET}
The analysis of the estimate computed here is based on standard techniques in analytic number theory.
\begin{lem} \label{lem2099ET.100B}\hypertarget{lem2099ET.100B} Let $x \geq 1$ be a large number and let $N\sim x$ be a large prime. Suppose there are no primes $p=n^2+1$ in the short interval $(x,2x]$. Then
\begin{equation} \label{eq2099.400d}
E(x)=\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \left (\frac{1}{N}\sum_{\sqrt{x}< m\leq \sqrt{2x},}\sum_{1\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }\right) = O\left ( (\log x)^{3}\right )\nonumber.
\end{equation}
\end{lem}
\begin{proof}[\textbf{Proof}] Partitioning the inner finite sum in the following way.
\begin{eqnarray} \label{eq2099.400g}
	E(x)	&=&\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \left (\frac{1}{N}\sum_{\sqrt{x}< m\leq \sqrt{2x},}\sum_{1\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }\right) \\[.3cm]
	&=&\frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{x}< m\leq \sqrt{2x},}\left (\sum_{1\leq u<N/2} e^{i2\pi \left(m^2-n \right)u/N }+\sum_{N/2\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }\right)  \nonumber\\[.3cm]
	&=&E_0(x) \; + \; E_1(x)\nonumber.
\end{eqnarray}
Summing the inner sum in the first subsum yields
\begin{eqnarray} \label{eq2099.400i}
	E_0(x)	&=&\frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1)\sum_{\sqrt{x}< m\leq \sqrt{2x},}\sum_{1\leq u<N/2} e^{i2\pi \left(m^2-n \right)u/N }\\[.3cm]
	&=& \frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{x}< m\leq \sqrt{2x},} \frac{e^{i2\pi (\frac{m^{2}-n}{N})(\frac{N+1}{2})}-1}{e^{i2\pi \frac{(m^{2}-n)}{N}}-1} \nonumber\\[.3cm]
	&\leq &\frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{x}< m\leq \sqrt{2x}}\left(\frac{2}{|\sin\pi(m^2-n)/N|} \right)	\nonumber.
\end{eqnarray}
The last inequality in \eqref{eq2099.400i} follows from the hypothesis $m^2-n\ne0$ and taking absolute value.\\

Replace $N\sim x$ and apply \hyperlink{lem2099SS.300S}{Lemma} \ref{lem2099SS.300S} to obtain the next inequality 
\begin{eqnarray} \label{eq2099.400t}
	E_{0}(x)&\leq &\frac{1}{N}\sum_{\substack{N< n\leq 2N\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{N}< m\leq \sqrt{2N}}\left(\frac{2}{|\sin\pi(m^2-n)/N|} \right)	\\[.3cm]
	&\ll&   \frac{\log 2N}{N}\sum_{a\leq \sqrt{2N},}  \sum_{b \leq N}\left(\frac{1}{|\sin\pi ab/N|} \right)	\nonumber ,
\end{eqnarray} 
where $|\sin\pi ab/N|\ne0$ since $N\nmid ab$. Observe that as the index in the first double sum in \eqref{eq2099.400t} ranges over $(m,n)\in \mathscr{M}\times (N,2N]$, it ranges over $\sqrt{2N}-\sqrt{N}-1$ copies of the set $(N,2N]-\{m^2\}$, where $\mathscr{M}=\{\sqrt{N}+1,\sqrt{N}+2,\ldots, \sqrt{2N}\}$. In contrast, as the index in the last double sum in \eqref{eq2099.400t} ranges over $(a,b)\in [1,\sqrt{2N}]\times [1,N]$ it ranges over $\sqrt{2N}$ copies of the set $[1,N]$. Thus, \hyperlink{lem2099SS.300S}{Lemma} \ref{lem2099SS.300S} fully justifies the inequality \eqref{eq2099.400t}. In light of this information, continuing the calculation of the estimate as in {\color{red}\cite[p.\;136]{DH2000}}, yields 

\begin{eqnarray} \label{eq2099.400u}
	E_0(x)	
	&\ll &\frac{\log N}{N}\sum_{a\leq \sqrt{2N},}  \sum_{b \leq N}\left(\frac{1}{|\sin\pi ab/N|} \right)	\\[.3cm]	 
	&\ll &\frac{\log N}{N}\sum_{a\leq \sqrt{2N},}  \sum_{b \leq N}\left(\frac{N}{\pi ab} \right)	\nonumber\\	 [.3cm] 
	&\ll &(\log N)\sum_{a\leq \sqrt{2N}}\frac{1}{a}  \sum_{b \leq N}\frac{1}{b}	\nonumber\\	[.3cm]  
	&\ll &(\log N)^3	\nonumber\\[.3cm]
		&\ll &(\log x)^3	\nonumber.
\end{eqnarray}
Similarly, the second subsum has the upper bound
\begin{eqnarray} \label{eq2099.400v}
	E_{1}(x)&=& \frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{m \leq \sqrt{x},}\sum_{N/2\leq u<N} e^{i2\pi \left(m^2-n \right)u/N } \\[.3cm]
	&=&    \frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{m \leq \sqrt{x},} \frac{1-e^{i2\pi (\frac{m^{2}-n}{N})(\frac{N+1}{2})}}{e^{i2\pi \frac{(m^{2}-n)}{N}}-1} \nonumber\\[.3cm]
	&\leq&   \frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{m \leq \sqrt{x}}\left(\frac{2}{|\sin\pi(m^2-n)/N|} \right) \nonumber\\[.3cm]
	&\ll&  (\log x)^3\nonumber.
\end{eqnarray}
Adding \eqref{eq2099.400u} and \eqref{eq2099.400v} and substituting \( x=(\log p)^{1+\varepsilon}\) yield
\begin{eqnarray} \label{eq2099.300x}
	E(x)&=& E_{0}(x)\;+\;E_{1}(x)  \\[.3cm]
	&\ll& (\log x)^3\nonumber.
\end{eqnarray}
This completes the verification.
\end{proof}

\section{Main Result} \label{S2099}\hypertarget{S2099}
The proof of the main result is based on the quadratic to linear identity described \hyperlink{S2070}{Section} \ref{S2070}.
\begin{proof}[\textbf{Proof}] (\hyperlink{thm9090.119}{Theorem} \ref{thm9090.119}) Take a sufficiently large number $x$ and suppose there are no primes $p=n^2+1\geq x$. Equivalently  $\Lambda(n^2+1)=0$ for all $n\geq x^{1/2}$. Summing the quadratic to linear identity in \hyperlink{lem2070.130}{Lemma} \ref{lem2070.130} over the short interval $(x,2x]$ leads to null sum
	\begin{equation} \label{eq2099.134d}
		\sum_{\sqrt{x}< m\leq \sqrt{2x}} \Lambda(m^2+1)= \sum_{\substack{x< n\leq 2x\\n\ne m^2}}\mathcal{I}_2(n)\Lambda(n+1)=0.
	\end{equation}
	This is a quantitative form of the hypothesis``\textit{there are no primes $p=n^2+1$ on the short interval} $(x,2x]$".\\
	
	Substituting the complete formula for the indicator function in \eqref{eq2070.130d} and expanding it yield 
	\begin{eqnarray} \label{eq2099.134f}
		\sum_{\sqrt{x}< m\leq \sqrt{2x}} \Lambda(m^2+1)
		&=& \sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \left (\frac{1}{N} \sum_{\sqrt{x}< m\leq \sqrt{2x},}\sum_{0\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }\right) \nonumber\\[.3cm]
		&=&\frac{1}{N}\sum_{\substack{x< n\leq 2x\\n\ne m^2}} \Lambda(n+1) \sum_{\sqrt{x}< m\leq \sqrt{2x}}1 \nonumber\\[.3cm]
		&&\hskip .0005 in +\sum_{\substack{x< n\leq 2x\\n\ne m^2}}\Lambda(n+1) \left (\frac{1}{N} \sum_{\sqrt{x}< m\leq \sqrt{2x},}\sum_{1\leq u<N} e^{i2\pi \left(m^2-n \right)u/N }\right)   \nonumber\\[.3cm]
		&=&M(x) \; + \; E(x).
	\end{eqnarray}
	The main term $M(x)$, which is associated with the value $u=0$, is computed in \hyperlink{lem2099MT.100A}{Lemma} \ref{lem2099MT.100A}. The error term, which is associated with the values $u=1,\ldots,N-1$ is estimated in \hyperlink{lem2099ET.100B}{Lemma} \ref{lem2099ET.100B}. Substituting these estimates yield
	\begin{eqnarray} \label{eq2099.134i}
		\sum_{\sqrt{x}< m\leq \sqrt{2x}} \Lambda(m^2+1)
		&=&M(x) \; + \; E(x)\\[.3cm]
		&\gg& \sqrt{x}\left( 1+O\left(e^{-c\sqrt{\log x}} \right) \right)+O\left( (\log x)^3\right) \nonumber\\[.3cm]
		&\gg& \sqrt{x}\left( 1+O\left(e^{-c\sqrt{\log x}} \right) \right) \nonumber\\[.3cm]
		&\gg& 1 \nonumber
	\end{eqnarray}
	for all large real numbers $x$, where $c>0$ is a constant. But this contradicts the hypothesis \eqref{eq2099.134d}. Therefore, the short interval $[x,2x]$ contains a prime $p=n^2+1$ for all large real numbers $x$.  
	Quod erat inveniendum. \end{proof}

The earliest numerical data seems to be the Euler table in {\color{red}\cite[p.\ 123]{EL1900}}. Some other numerical data and experiments are reported in {\color{red}\cite[p.\ 50]{HL1923}}, \cite{SD1959}, and \cite{RI2015}. An instructive numerical experiment for the polynomial $f(x)=x^2+1$ is conducted in {\color{red}\cite[Section 3.3]{FG2018}}.



\end{document}